OpenStudy (anonymous):

Gabrielle is 5 years older than Mikhail. The sum of their ages is 49. What is Mikhail's age?

eclipsedstar (eclipsedstar):

Sum is another word for Total.

eclipsedstar (eclipsedstar):

@Maria195 I don't think one could have 29.5 as an age....just saying XD

OpenStudy (anonymous):

Have you ever thought of 291/2? :P

OpenStudy (anonymous):

29 and a half

eclipsedstar (eclipsedstar):

Lol,

OpenStudy (anonymous):

Gahh, this is confusing lol

eclipsedstar (eclipsedstar):

I thought it was a trick question too

OpenStudy (anonymous):

yeah but it doesn't say she's 5, it says she's 5 years older

eclipsedstar (eclipsedstar):

Ik I just saw that XD

eclipsedstar (eclipsedstar):

well, @Maria195 your first reply on the very top doesn't add up...

OpenStudy (anonymous):

I'll just delete it then

OpenStudy (anonymous):

49 - 5 [years old then Mikhail's] = 44 44 / 2 = 22 22 {Mikhail's age) + 22 {+5} = 27 27 + 22 = 49

OpenStudy (anonymous):

So Mikhail is 22 years old

eclipsedstar (eclipsedstar):

Okay that makes sense. ^

OpenStudy (anonymous):

Well, I guess we all benefited from that XD

OpenStudy (anonymous):

thanks